OpenStudy (anonymous):

(2y-2)^3 assume y = 0

OpenStudy (anonymous):

i answerd it dear now if we put y=0 in equation we will get -8

OpenStudy (anonymous):

lizzy, if y = 0 then the equation becomes (2(0) - 2)^3 = (0-2)^3 = -2^3 = -8

OpenStudy (anonymous):

that is not a choice answer.. they give u multiple choice for this problem??? so it must not be correct.. sorry..

OpenStudy (anonymous):

well, it doesnt matter. the answer is -8 whether its a choice or not. maybe your book is wrong.

OpenStudy (anonymous):

or it has a printing mistake.

OpenStudy (anonymous):

I think you may have given the question wrong. I don't see why you would 'assume' y = 0

OpenStudy (anonymous):

thanks for trying to help